OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (–2, –3) and is parallel to the line y = x + 9?

OpenStudy (anonymous):

x – y = 1 x + y = 1 x – y = –5 x + y = 5 i chose the last one

OpenStudy (anonymous):

parallel mean that the new line have the same slope of you line: so it will have equation => y=x+q sostitute the x and y of the point, you'll have -3=-2+q => q=-1 so the solution is y=x-1 , the first one!